# MinatoLoader: Accelerating Machine Learning Training Through Efficient Data Preprocessing

### v1

@@ Abstract.
Data loaders are used by Machine Learning (ML) frameworks like PyTorch and TensorFlow to apply transformations to data before feeding it into the accelerator. This operation is called data preprocessing. Data preprocessing plays an important role in the ML training workflow because if it is inefficiently pipelined with the training, it can yield high GPU idleness, resulting in important training delays. Unfortunately, existing data loaders turn out to waste GPU resources, with $76\%$ GPU idleness when using the PyTorch data loader, for example. One key source of inefficiency is the variability in preprocessing time across samples within the same dataset. Existing data loaders are oblivious to this variability, and they construct batches without any consideration of slow or fast samples. In this case, the entire batch is delayed by a single slow sample, stalling the training pipeline and resulting in head-of-line blocking.
To address these inefficiencies, we present MinatoLoader, a general-purpose data loader for PyTorch that accelerates training and improves GPU utilization. MinatoLoader is designed for a single-server setup, containing multiple GPUs. It continuously prepares data in the background and actively constructs batches by prioritizing fast-to-preprocess samples, while slower samples are processed in parallel.
We evaluate MinatoLoader on servers with V100 and A100 GPUs. On a machine with four A100 GPUs, MinatoLoader improves the training time of a wide range of workloads by up to $7.5\times$ ($3.6\times$ on average) over PyTorch DataLoader and Pecan, and up to $3\times$ ($2.2\times$ on average) over DALI. It also increases average GPU utilization from 46.4% with PyTorch to 90.45%, while preserving model accuracy and enabling faster convergence.
@@ 1. Introduction
The efficacy of Machine Learning (ML) deployments relies on both high-quality algorithms and high-quality data, the latter being shaped through data preprocessing. Although considerable research focused on improving the training algorithms – leading to advances in techniques (Kotsiantis et al., 2006; Maetschke et al., 2017; Reddy et al., 2020), software libraries (num, [n. d.]; sci, [n. d.]; pan, [n. d.]; Mattson et al., 2020), and hardware accelerators (e.g., GPUs, TPUs, DPUs)– the data preprocessing step has been relatively underexplored (Gao et al., 2024; Sambasivan et al., 2021). Yet, data preprocessing plays an important role in training efficiency and model generalization. Transformations such as cropping, resizing, shuffling, and random augmentations are commonly applied on the fly during training to increase model robustness and accuracy (Zhao et al., 2022b; Gupta et al., 2022; Mazumder et al., 2022; Shorten and Khoshgoftaar, 2019). These operations expose the model to diverse inputs, improving generalization and training speed.
@@
(b) CPU and GPU usage of PyTorch DataLoader during 3D-UNet training.
@@
In this paper, we address the following question: How can the data loader deliver preprocessed data to the GPU as fast as possible to minimize idle periods? To achieve this, we propose MinatoLoader, a general-purpose, drop-in replacement for the PyTorch DataLoader that improves training time and GPU utilization without requiring prior knowledge of the dataset or preprocessing pipeline. This work extends our previous workshop paper (Nouaji, [n. d.]). Contrary to prior work (pyt, 2025; Graur et al., 2024; dal, [n. d.]), which treats all data samples uniformly and blocks batch creation on the slowest sample, MinatoLoader introduces a sample-aware scheduling strategy that dynamically adapts to per-sample preprocessing variability. The key idea behind MinatoLoader is to prioritize fast-to-process samples and defer slow ones for later use. This is a simple yet powerful approach that improves training performance, is generalizable across workloads, and maintains virtually the same model accuracy despite sample reordering inside the batches.
@@
  • We combine the above-mentioned observations and techniques into MinatoLoader, a general-purpose data loader that proactively prepares training data without requiring any prior knowledge of the dataset or preprocessing pipeline. We open source MinatoLoader at https://anonymous.4open.science/r/eurosys-submission-1683/.
@@ 2.1. Data Loaders
PyTorch DataLoader (pyt, 2025) uses a standard iterator to access the dataset. It allows parallel data loading with worker threads, where each worker can prefetch several samples (controlled by prefetch_factor). While these settings help increase throughput, they are limited when preprocessing becomes the bottleneck. This limitation is rooted in how PyTorch constructs and processes batches. It first defines the order in which data indices are drawn from the dataset, and then groups these indices into batches. However, by default, the data loader is unaware of preprocessing cost differences between samples. Consequently, a batch may contain both fast and slow samples, leading to head-of-line blocking where an entire batch is delayed by its slowest sample, as illustrated in Figure 1(a).
@@ 3. Issues in the Data Preprocessing Pipeline
We use three representative workloads from the MLPerf Training Benchmark, using different models and preprocessing pipelines, as depicted in Table 1 and previously described in section˜2.2.
@@ 3.1. Preprocessing Time Variability
Takeaway 1.
(a) Image segmentation (3D-UNet).
(b) Object detection (R-CNN).
@@ 3.4. Tuning Prefetching Parameters
Figure 4 depicts the overall training time under different configurations. While both mechanisms aim to overlap data loading with training, results show that increasing their values yields limited benefits because neither system reduces the per-sample transformation cost. In the PyTorch DataLoader, increasing the prefetch factor offers minimal (even almost nil) improvement and may cause out-of-memory errors due to excessive buffering. In addition, in DALI, where preprocessing is offloaded to the GPU, higher prefetch queue depth increases GPU memory usage and preprocessing load, which can interfere with training computations and ultimately prolong overall training time.
@@ 4. Design and Implementation
MinatoLoader is a general-purpose data loader that proactively prepares training data to enable efficient batch construction, without requiring any prior knowledge of the dataset or preprocessing pipeline. MinatoLoader is designed for local training scenarios (i.e., single node, multiple GPUs) and is provided as a drop-in replacement of PyTorch DataLoader, following the same API. Next, we present an overview of the MinatoLoader architecture (section˜4.1), detail the key techniques that underpin its design (section˜4.2-section˜4.3), and discuss implementation details (section˜4.4).
@@ 4.1. MinatoLoader in a Nutshell
Figure 5. MinatoLoader high-level design. It continuously enqueues preprocessed samples into a specified queue based on a load balancer decision. Concurrently, the GPU dequeues preprocessed data for training. We show MinatoLoader for one GPU, but it can generalize to multi-GPU settings.
Figure 5 depicts the high-level architecture of MinatoLoader, which is built around three key stages designed to prevent pipeline stalls and maximize GPU utilization: data loading, preprocessing, and batch construction. For clarity, we show how MinatoLoader operates for one GPU, but it can generalize to multi-GPU settings, as shown in our experiments (section˜5). First, the data loading stage is identical to PyTorch DataLoader, following the same API. Second, at the core of the system is a dynamic, sample-aware load balancer (section˜4.2) designed for mitigating head-of-line blocking caused by slow preprocessing samples (section˜3) by classifying samples as fast or slow on-the-fly during training. To achieve this, it maintains four queues: a fast queue for quick preprocessing samples, a slow queue for samples exceeding a predefined preprocessing timeout, a batch queue that assembles training batches using both fast and slow queues, and a temp queue used to temporarily store long samples that need to be preprocessed off the critical path. MinatoLoader maintains one batch queue per GPU, and separate fast, slow, and temp queues for each CPU worker. These queues enable decoupling data preparation from batch construction, allowing MinatoLoader to prioritize readily available samples and delay slower ones without stalling the pipeline.
Further, MinatoLoader includes a worker scheduler (section˜4.3) that dynamically tunes the number of CPU preprocessing workers to match the throughput demands of the training workers on the GPUs (omitted in Figure 5 for clarity).
@@ 4.2. Avoiding Head-of-Line Blocking
Algorithm 1 illustrates the core logic of MinatoLoader’s load balancer. Given a sample $s$ and a list of transformations $T$, MinatoLoader applies the transformations sequentially while monitoring the elapsed time (7–10). If all transformations are completed within the timeout budget $t_{out}$, the sample is enqueued into the fast_queue (12). If the timeout $t_{out}$ is exceeded, the system interrupts preprocessing, records the index of the transformation that was in progress, and enqueues it along with its index into the temp_queue (11). Since the last transformation was only partially applied, it must be re-executed to ensure that the sample is in a valid state to carry on the remaining transformations. This design avoids restarting the entire data preprocessing pipeline.
Afterwards, a background worker resumes the preprocessing of the slow samples from the recorded index and, once complete, places them into the slow_queue (14–18). Meanwhile, a batch construction thread continuously builds batches by pulling from both fast_queue and slow_queue (22–30). Finally, MinatoLoader continuously dequeues ready-to-train batches from the batch_queue (32–37). Empirically, we determine 10 ms as the optimal sleep time (28 and 37).
@@ 4.3. Keeping the GPUs Busy
The CPU workers include the number of data loading workers, the slow-task workers, and the batch workers. The data loading CPU workers are responsible for loading data from the disk, preprocessing the fast samples, and moving the slow samples to the temp queues. Slow-task CPU workers run in the background, without blocking the pipeline. The batch CPU workers pull samples from the slow and fast queues to construct batches ahead of time and ensure that the GPU workers always have prebuilt batches to ingest. In addition, MinatoLoader incorporates a prefetching mechanism that uses a CUDA stream to transfer the batch $i$ from the batch queue to the GPU memory before the GPU has finished executing the batch i1i-1, which helps achieve continuous GPU training.
@@ 5.5. Performance Under Memory Constraints
Figure 10. CPU and GPU usage (left) and disk read (right) of the image segmentation workload with PyTorch DataLoader ($\approx$650s), DALI ($\approx$500s), and MinatoLoader ($\approx$330s) when training a 230GB dataset under an 80GB memory constraint.
@@
5.6. Accuracy Preserving with MinatoLoader
(a) Object detection (R-CNN)
(b) Image segmentation (3D-UNet)
Figure 11. Accuracy of (a) Mask R-CNN and (b) 3D-UNet with PyTorch DataLoader and MinatoLoader. Since Speech is a microbenchmark, we did not evaluate its accuracy here.
@@ 5.6. Accuracy Preserving with MinatoLoader
For image segmentation, we trained the model until accuracy stabilized, which required approximately 500 epochs. This took 8 hours and 2 minutes with PyTorch DataLoader and only 3 hours and 52 minutes with MinatoLoader. As illustrated in Figure 11(b), MinatoLoader preserves the full accuracy trend and reaches the same final accuracy of 58%.
@@ 6. Related Work
Numerous systems aim to improve ML training efficiency by optimizing various stages of the data pipeline. Some focus on the input pipeline (Graur et al., 2024; Murray et al., 2021; Um et al., 2023; Kuchnik et al., 2019; Zhao et al., 2022a, 2023a), while others cache preprocessed data (Graur et al., 2022; Mohan et al., 2021; Lee et al., 2021). Some offload preprocessing to GPUs or FPGAs (dal, [n. d.]; Kim et al., 2023; Park et al., 2020), and others use storage-centric solutions to accelerate data preprocessing (Lee et al., 2024; Zhao et al., 2023b). Finally, there is work on tools to characterize the performance of data preprocessing for ML (Bachkaniwala et al., 2024; Balmau, 2022).

### v2

@@ Abstract.
Machine learning (ML) frameworks, such as PyTorch and TensorFlow, rely on data loaders to preprocess data before feeding it to accelerators. When preprocessing is inefficiently pipelined, GPUs can remain idle over long periods of time, leading to substantial training delays. For example, PyTorch’s default data loaders can cause up to 76% GPU idleness. A key bottleneck is the variability in preprocessing time across samples within the same dataset. Existing data loaders are oblivious to this variability, training all samples uniformly. In this case, a single slow sample can stall the entire batch, causing head-of-line blocking.
We present MinatoLoader, a general-purpose data loader for PyTorch that accelerates training and improves GPU utilization under single-server, multi-GPU settings. It continuously prepares data in background and constructs batches by prioritizing fast-to-process samples, while slower samples are processed in parallel.
[…]
ML training, Data loader, Data preprocessing
journalyear: 2026conference: European Conference on Computer Systems; April 27–30, 2026; Edinburgh, Scotland Ukbooktitle: European Conference on Computer Systems (EUROSYS ’26), April 27–30, 2026, Edinburgh, Scotland Ukdoi: 10.1145/3767295.3769376isbn: 979-8-4007-2212-7/26/04ccs: Computing methodologies Machine learningccs: Computer systems organization Data flow architecturesccs: Computing methodologies Parallel algorithms
@@ 1. Introduction
The efficacy of Machine Learning (ML) deployments relies on both high-quality algorithms and high-quality data, the latter being shaped through data preprocessing. Although considerable research focused on improving the training algorithms – leading to advances in techniques (Kotsiantis et al., 2006; Maetschke et al., 2017; Reddy et al., 2020), software libraries (num, [n. d.]; sci, [n. d.]; pan, [n. d.]; Mattson et al., 2020), and hardware accelerators (e.g., GPUs, TPUs, DPUs)– the data preprocessing step has been relatively underexplored (Gao et al., 2024; Sambasivan et al., 2021; Nouaji et al., 2024). Yet, data preprocessing plays an important role in training efficiency and model generalization. Transformations such as cropping, resizing, shuffling, and random augmentations are commonly applied on the fly during training to increase model robustness and accuracy (Zhao et al., 2022; Gupta et al., 2022; Mazumder et al., 2023; Shorten and Khoshgoftaar, 2019). These operations expose the model to diverse inputs, improving generalization and training speed.
@@
(b) CPU and GPU usage of PyTorch DataLoader during 3D‑UNet training.
@@
In this paper, we address the following question: How can the data loader deliver preprocessed data to the GPU as fast as possible to minimize idle periods?
To achieve this, we propose MinatoLoader, a general-purpose, drop-in replacement for the PyTorch DataLoader that improves training time and GPU utilization without requiring prior knowledge of the dataset or preprocessing pipeline. Contrary to prior work (pyt, 2025; Graur et al., 2024; dal, [n. d.]), which treats all data samples uniformly and blocks batch creation on the slowest sample, MinatoLoader introduces a sample-aware scheduling strategy that dynamically adapts to per-sample preprocessing variability. The key idea behind MinatoLoader is to prioritize fast-to-process samples and defer slow ones for later use. This is a simple yet powerful approach that improves training performance, is generalizable across workloads, and maintains virtually the same model accuracy despite sample reordering inside the batches.
@@
  • We combine the above-mentioned observations and techniques into MinatoLoader, a general-purpose data loader that proactively prepares training data without requiring any prior knowledge of the dataset or preprocessing pipeline. We open source MinatoLoader at https://github.com/Rahm-no/MinatoLoader and Zenodo (Nouaji and Bitchebe, 2025).
@@ 2.1. Data Loaders
PyTorch DataLoader (pyt, 2025) uses a standard iterator to access the dataset and supports parallel data loading with worker threads, where each worker can prefetch several samples (controlled by prefetch_factor). While these settings help increase throughput, they are limited when preprocessing becomes the bottleneck. The issue lies in how PyTorch constructs and processes batches. It first defines the order in which data indices are drawn from the dataset, and then groups these into batches. However, by default, the data loader is unaware of differences in preprocessing cost. Consequently, a batch may contain both fast and slow samples, leading to head-of-line blocking where an entire batch is delayed by its slowest sample, as illustrated in Figure 1(a).
@@ 3. Issues in the Data Preprocessing Pipeline
We use three representative workloads from the MLPerf Training Benchmark, with different models and preprocessing pipelines, as described in section˜2.2 and Table 1.
@@ 3.4. Tuning Prefetching Parameters
Figure 4 depicts the overall training time under different configurations. While both mechanisms aim to overlap data loading with training, results show that increasing their values yields limited benefits because neither system reduces the per-sample transformation cost. In the PyTorch DataLoader, increasing the prefetch factor offers minimal (even almost nil) improvement and may cause out-of-memory (OOM) errors due to excessive buffering. In addition, in DALI, where preprocessing is offloaded to the GPU, higher prefetch queue depth increases GPU memory usage and preprocessing load, which can interfere with training computations and ultimately prolong overall training time.
@@ 4. Design and Implementation
MinatoLoader is a general-purpose data loader that proactively prepares training data to enable efficient batch construction, without requiring prior knowledge of the dataset or preprocessing pipeline. It is designed for local training scenarios (i.e., single node, multiple GPUs) and is provided as a drop-in replacement of PyTorch DataLoader, following the same API. MinatoLoader also supports distributed training (section˜6). Next, we present an overview of the MinatoLoader architecture (section˜4.1), detail the key techniques that underpin its design (section˜4.2-section˜4.3), and discuss implementation details (section˜4.4).
@@ 4.1. MinatoLoader in a Nutshell
Figure 5. MinatoLoader high-level design. It continuously enqueues preprocessed samples into a specified queue based on a load balancer decision. Concurrently, the GPU dequeues preprocessed data for training. We show MinatoLoader for one GPU, but it generalizes to multi-GPU settings.
Figure 5 depicts the high-level architecture of MinatoLoader, which is built around three key stages designed to prevent pipeline stalls and maximize GPU utilization: data loading, preprocessing, and batch construction. For clarity, we show how MinatoLoader operates for one GPU, but it generalizes to multi-GPU settings, as shown in our experiments (section˜5). First, the data loading stage is identical to PyTorch DataLoader, following the same API. Second, at the core of the system is a dynamic, sample-aware load balancer (section˜4.2) designed for mitigating head-of-line blocking caused by slow preprocessing samples (section˜3) by classifying samples as fast or slow on-the-fly during training. To achieve this, it maintains four queues: a fast queue for quick preprocessing samples, a slow queue for samples exceeding a predefined preprocessing timeout, a batch queue that assembles training batches using both fast and slow queues, and a temp queue used to temporarily store long samples that need to be preprocessed off the critical path. MinatoLoader maintains one batch queue per GPU, and separate fast, slow, and temp queues for each CPU worker. These queues enable decoupling data preparation from batch construction, allowing MinatoLoader to prioritize readily available samples and delay slower ones without stalling the pipeline. Further, MinatoLoader includes a worker scheduler (section˜4.3) that dynamically tunes the number of CPU preprocessing workers to match the throughput demands of the training workers on the GPUs (omitted in Figure 5 for clarity).
Like PyTorch, MinatoLoader requests samples in random order, but it builds batches from whichever samples finish preprocessing first, rather than imposing a strict order. Additionally, slow samples are never preempted: they continue preprocessing in background and are added to future batches as soon as they are ready. Importantly, MinatoLoader does not defer these samples to the very end, ensuring that its randomized batches remain statistically similar to PyTorch Dataloader, as shown in section˜5.6.
@@ 45 MinatoLoader __next__() Call
Algorithm 1 illustrates the core logic of MinatoLoader’s load balancer. Given a sample $s$ and a list of transformations $T$, MinatoLoader applies the transformations sequentially while monitoring the elapsed time (7-10). If all transformations are completed within the timeout budget $t_{out}$, the sample is enqueued into the fast_queue (12). If the timeout $t_{out}$ is exceeded, the system interrupts preprocessing, records the index of the transformation that was in progress, and enqueues it along with its index into the temp_queue (11). Since the last transformation was only partially applied, it must be re-executed to ensure that the sample is in a valid state to carry on the remaining transformations. This design avoids restarting the entire data preprocessing pipeline.
Afterwards, a background worker resumes the preprocessing of the slow samples from the recorded index and, once complete, places them into the slow_queue (14-18). Meanwhile, a batch construction thread continuously builds batches by pulling from both fast_queue and slow_queue (22-30). Finally, MinatoLoader continuously dequeues ready-to-train batches from the batch_queue (32-37). Empirically, we determine 10 ms as the optimal sleep time (28 and 37).
@@ 4.3. Keeping the GPUs Busy
The CPU workers include the number of data loading workers, the slow-task workers, and the batch workers. The data loading CPU workers are responsible for loading data from disk, preprocessing fast samples, and moving slow samples to the temp queues. Slow-task CPU workers run in the background, without blocking the pipeline. The batch CPU workers pull samples from the slow and fast queues to construct batches ahead of time and ensure that the GPU workers always have prebuilt batches to ingest. In addition, MinatoLoader incorporates a prefetching mechanism that uses a CUDA stream to transfer the batch $i$ from the batch queue to the GPU memory before the GPU has finished executing the batch i1i-1, which helps achieve continuous GPU training.
@@ 5. Evaluation
  • Q5: How does MinatoLoader behave under varying proportions of slow samples (section˜5.6)?
@@ 5.5. Performance Under Memory Constraints
Figure 10. CPU and GPU usage (left) and disk read (right) of the image segmentation workload with PyTorch DataLoader ($\approx$650s), DALI ($\approx$500s), and MinatoLoader ($\approx$330s) when training a 230GB dataset under an 80GB memory limit.
@@
5.6. Sensitivity Analysis
We now evaluate distinct features of MinatoLoader.
(a) Accuracy of PyTorch DataLoader and MinatoLoader.
(b) Distribution of batches by the number of slow samples.
(c) Proportion of slow samples over training iterations. Dashed lines show the average fraction of slow samples. For readability, only a subset of the execution is shown.
Figure 12. Sensitivity analysis of PyTorch DataLoader (in blue) and MinatoLoader (in orange) with object detection (left) and image segmentation (right) workloads.
Accuracy Preserving with MinatoLoader
@@ 5.6. Sensitivity Analysis
For image segmentation, we trained the model until accuracy stabilized, which required approximately 500 epochs. This took 8 hours and 2 minutes with PyTorch DataLoader and only 3 hours and 52 minutes with MinatoLoader. As illustrated in Figure 12(a), MinatoLoader preserves the full accuracy trend and reaches the same final accuracy of 58%.
Batch composition analysis
To complement the accuracy results, we conducted a batch composition study to validate that MinatoLoader’s strategy does not introduce bias. Specifically, we evaluated (i) the distribution of batches by the number of slow samples they contain (Figure 12(b)), i.e., the proportion of batches with a given number of samples, and (ii) the proportion of slow samples over training iterations (Figure 12(c)). Experiments were conducted on the object detection (R-CNN) and image segmentation (3D-UNet) workloads with a batch size of 4.
Results show that MinatoLoader constructs mixed batches comparable to PyTorch DataLoader and does not introduce systematic bias. Figure 12(b) shows a similar distribution of batches across systems for both workloads, indicating that MinatoLoader preserves the natural ratio of slow samples. Figure 12(c) further shows that slow samples are incorporated into batches as soon as they are ready, rather than being deferred or preempted. For example, PyTorch DataLoader yields an average proportion of slow samples of 0.15 and 0.23 under the object detection and image segmentation workloads, respectively, while MinatoLoader presents 0.17 and 0.24. These results confirm that MinatoLoader preserves fairness in sample usage and maintains a batch distribution closely aligned with PyTorch DataLoader.
Cluster of slow samples
We now evaluate MinatoLoader under varying proportions of slow and fast samples. To this end, we modified the Speech-3s workload so that the HeavyStep transformation is applied to a configurable proportion of the dataset, rather than every five samples. Experiments were conducted with increasing proportion of slow samples from 0% up to 100%. Figure 13 depicts the training time of PyTorch DataLoader, DALI, Pecan, and MinatoLoader. As expected, in the edge cases (i.e., 0% and 100%), MinatoLoader performs similarly to PyTorch DataLoader and Pecan, since all samples have uniform preprocessing cost. The benefits of MinatoLoader appear in the intermediate range (25%–75%), whose strategy leverages the variability across data samples, outperforming existing solutions up to $2.4\times$.
Figure 13. Training time of PyTorch DataLoader, Pecan, DALI, and MinatoLoader across different proportions of slow samples.
6. Discussion
Order-sensitive scenarios
Some training scenarios require strict sample ordering. In multimodal models (e.g., image-text or audio-text), it is essential to maintain alignment between paired modalities. In the speech recognition workload (section˜2.2), MinatoLoader orders samples based on audio preprocessing time but always processes the audio–text pair together, preventing mismatches. This approach generalizes to other multimodal settings. Similarly, in curriculum learning (Soviany et al., 2022), where the global sample order is semantically important (e.g., easier examples preceding harder ones), MinatoLoader can be configured to disable reordering. In this mode, it behaves like PyTorch DataLoader, preserving the required sample order. While this disables MinatoLoader ’s reordering advantage, it guarantees correctness in scenarios that depend on strict order.
Distributed training
While our evaluation focused on a single server, MinatoLoader’s design generalizes for distributed training with multiple nodes and GPUs. In such scenarios, each server has a PyTorch instance accessing data from local or shared storage, with MinatoLoader integrated into each instance. Under data and model parallelism, MinatoLoader retains its preprocessing and batch construction benefits, and performs similarly to PyTorch’s DataLoader during the data loading phase, as shown in Figure 5.
Suitable workloads
MinatoLoader provides the largest benefits for workloads where preprocessing is heavy, such as image, speech, or video training pipelines. In contrast, text-based NLP workloads, including large language models, are typically not input-bound, as their preprocessing is lightweight (e.g., tokenization) or performed offline. In these cases, MinatoLoader offers limited additional benefit and behaves similarly to the default PyTorch DataLoader. A key future challenge for efficient ML preprocessing lies in multimodal workloads (Liang et al., 2024), where different data modalities impose distinct preprocessing requirements, making this stage critical for training performance.
@@ 7. Related Work
Numerous systems aim to improve ML training efficiency by optimizing various stages of the data pipeline. Some focus on the input pipeline (Graur et al., 2024; Murray et al., 2021; Um et al., 2023; Kuchnik et al., 2021; Zhao et al., 2022, 2023a), while others cache preprocessed data (Graur et al., 2022; Mohan et al., 2021; Lee et al., 2021). Some offload preprocessing to GPUs or FPGAs (dal, [n. d.]; Kim et al., 2023; Park et al., 2020), and others use storage-centric solutions to accelerate data preprocessing (Lee et al., 2024; Zhao et al., 2023b). Finally, there is work on tools to characterize the performance of data preprocessing for ML (Bachkaniwala et al., 2024; Balmau, 2022), as well as big data systems techniques that address data preprocessing (Spark, 2019).
@@
Big data systems
MinatoLoader targets the preprocessing computation and batch construction phases, whereas techniques applied by traditional big data systems (e.g., data skew/straggler, dynamic repartitioning, workload estimation) target how data is organized, being complementary. Speculative execution is not applicable in ML preprocessing. Adaptive scheduling is the closest technique to MinatoLoader. However, while ML preprocessing has simpler dependencies between samples and workers, it has stricter latency requirements than adaptive scheduling. Prior work (Murray et al., 2021) also notes this difference. For example, Spark Streaming (spa, 2025) recommends batch granularities of at least 50ms, while ML training requires step times below 1ms. Further, in production, Spark is used earlier in the pipeline for ETL tasks (Zhao et al., 2022), with outputs written to storage and later consumed by PyTorch. This separation of concerns reflects current design choices: Spark addresses offline preprocessing, while PyTorch performs online transformations.